OpenStudy (anonymous):

can someone please show me how to use the convolution theorem to solve for the inverse laplace transform of s/(s^2+a^2)(s^2+b^2)

OpenStudy (anonymous):

First, you do the following: \[\mathcal{L}_s^{-1}\left[\frac{1}{a^2+s^2}\right](t)=\frac{\sin (a t)}{a}\\ \mathcal{L}_s^{-1}\left[\frac{s}{b^2+s^2}\right](t)=\ cos (b t) \]

OpenStudy (anonymous):

Now you can use that the Inverse Laplace transform of the product is the convolution of their inverse transform

OpenStudy (anonymous):

Then you do the work to find that \[ \mathcal{L}_s^{-1}\left[\frac{s}{\left(a^2+s^2\right) \left(b^2+s^2\right)}\right](t)=\int_0^t \frac{\sin (a s) \cos (b (t-s))}{a} \, ds=\frac{\cos (b t)-\cos (a t)}{a^2-b^2} \]

OpenStudy (anonymous):

\[ \mathcal{L}_s^{-1}\left[\frac{s}{\left(a^2+s^2\right) \left(b^2+s^2\right)}\right](t)=\\\int_0^t \frac{\sin (a s) \cos (b (t-s))}{a} \, ds=\frac{\cos (b t)-\cos (a t)}{a^2-b^2} \]

OpenStudy (anonymous):

thank you so so so much!!!

OpenStudy (anonymous):

YW